OpenStudy (anonymous):

Let f(x) = 12 over the quantity of 4 x + 2. Find f(-1).

OpenStudy (anonymous):

-6 -2 2 6

OpenStudy (anonymous):

@kirbykirby @wyattp17 @Jacob_Junior

OpenStudy (anonymous):

@PaulaLovesSchool13

OpenStudy (anonymous):

.....

OpenStudy (anonymous):

Reminder: I am not the best in math XD

hartnn (hartnn):

to get f(-1), just plug in x=-1 in f(x)!

OpenStudy (anonymous):

is it -6

OpenStudy (anonymous):

or 6

hartnn (hartnn):

\(\large f(x) = \dfrac{12}{4x+2} \\ \Large f(-1) = \dfrac{12}{4(-1)+2} = ... ?\)

hartnn (hartnn):

yes! -6 is correct :)

OpenStudy (anonymous):

Let f(x) = 2x - 6. Solve f-1(x) when x = 2. -5 4 -2 7

OpenStudy (anonymous):

is it b

hartnn (hartnn):

2x-6 = 2 YES! answer is 4, good! :)

OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). -7 -3 10 33

OpenStudy (anonymous):

here i dont know what to do

hartnn (hartnn):

find g(2) first, as usual

hartnn (hartnn):

plug in x=2 in g(x) to get g(2)

OpenStudy (anonymous):

-15 + 8

hartnn (hartnn):

excellent work!

hartnn (hartnn):

-7 is correct :)

OpenStudy (anonymous):

can u help me with two more

OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)). -21 -3 3 21

hartnn (hartnn):

sure :)

hartnn (hartnn):

whats g(-1) ? plug in x=-1 in g(x)

OpenStudy (anonymous):

Let f(x) = x2 - 81. Find f-1(x). ±square root of the quantity x plus 81 9square root of x 1 over quantity x squared minus 81 x squared over 81

OpenStudy (anonymous):

the llast oone i think its c

OpenStudy (anonymous):

and the 1 -21

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

good explaining Taylor! Wear this medal :)

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

f(g(-1)) is actually -3. how u got -21 ?

OpenStudy (anonymous):

i just did 7times 3

hartnn (hartnn):

g(-1) = 1 - 7(-1) - 9 = 1+7-9 = -1 f(x) = x - 2 , f(-1) = -1-2 =-3

hartnn (hartnn):

y = x^2 - 81 try to isolate x from here, that will be your inverse function

OpenStudy (anonymous):

- 81 times 2

hartnn (hartnn):

no...like this : \(y = x^2 -81 \\ y+81 =x^2 \\x = \pm \sqrt {y+81} \)

hartnn (hartnn):

so your inverse function will be square root of (x+81)

hartnn (hartnn):

"±square root of the quantity x plus 81"